# Caption This, Reason That: VLMs Caught in the Middle

### v1

@@ Abstract
Vision-Language Models (VLMs) have shown remarkable progress in visual understanding in recent years. Yet, they still lag behind human capabilities in specific visual tasks such as counting or relational reasoning. To understand the underlying limitations, we adopt methodologies from cognitive science, analyzing VLM performance along core cognitive axes: Perception, Attention, and Memory. Using a suite of tasks targeting these abilities, we evaluate state-of-the-art VLMs, including GPT-4o. Our analysis reveals distinct cognitive profiles: while advanced models approach ceiling performance on some tasks (e.g. category identification), a significant gap persists, particularly in tasks requiring spatial understanding or selective attention. Investigating the source of these failures and potential methods for improvement, we employ a vision-text decoupling analysis, finding that models struggling with direct visual reasoning show marked improvement when reasoning over their own generated text captions. These experiments reveal a strong need for improved VLM Chain-of-Thought (CoT) abilities, even in models that consistently exceed human performance. Furthermore, we demonstrate the potential of targeted fine-tuning on composite visual reasoning tasks and show that fine-tuning smaller VLMs substantially improves core cognitive abilities. While this improvement does not translate to large enhancements on challenging, out-of-distribution benchmarks, we show broadly that VLM performance on our datasets strongly correlates with performance on these other benchmarks. Our work provides a detailed analysis of VLM cognitive strengths and weaknesses and identifies key bottlenecks in simultaneous perception and reasoning while also providing an effective and simple solution.
@@ 2 Related Works
Beyond these broad vision-language benchmarks, recent studies have focused on testing VLMs on more narrow cognitive domains. BlindTest [48] measures the abilities of VLMs to solve visual geometric tasks. VCog-Bench [49] takes inspiration from early neuro-developmental cognitive tests and evaluates VLMs on their ability to solve abstract logic puzzles and pattern recognition tasks. VisFactor from [50] measures similar highly abstract visual reasoning abilities. Schulze Buschoff et al. [51] focus their cognitive evaluations on intuitive physics, causal reasoning, and intuitive psychology. However, by far the most commonly measured cognitive domain is spatial reasoning [28, 29, 30, 23, 31].
@@ 3.1 Data Source
We utilize the iWISDM task environment [22] to generate all cognitive tasks and fine-tuning data in this study. This environment enables the procedural generation of an effectively limitless number of vision-language decision-making tasks. In this study, we leverage iWISDM to generate tasks with varying levels of complexity, aligning with the requirements of each cognitive axis in the PAM dataset. These tasks range from simple single-object localization to more complex ones requiring logical reasoning and object comparisons across image sequences.
@@ 4.1 Cognitive Evaluation of VLMs with PAM Dataset
[…]
@@ 4.2.1 Relationship Between PAM & CVR
Intuitively, models with stronger memory, attention, and perception abilities should also perform better on tasks that require a combination of these skills. Since our CVR tasks are designed to engage exactly these abilities in combination, we plot each model’s PAM performance against its CVR performance, as shown in Figure 2. All performances on all three core cognitive axes showed strong correlation to CVR task performance. Memory task performance stands out with a correlation of 0.94. This is likely due to the abundance of Memory comparison-based subtasks (i.e comparing two objects with irrelevant images in between), contained within CVR tasks. Furthermore, models PAM and CVR accuracies are significantly correlated to their performance on widely used benchmarks such as MMMU-Pro (Figures A.6.1), which further validates the effectiveness of the evaluations presented here.
@@ 4.3.1 Decoupling Vision & Text via Captioning
[…]
@@
Perception (Loc) 44.33±5.59 +28.67±5.00 +10.17±6.84 +55.00±1.11
Feature Attention 55.85±2.65 +22.00±2.21 -1.74±3.25 +22.22±2.21
Spatial Attention 57.48±2.63 +18.15±2.29 -1.70±3.24 +8.52±2.52
Memory (Cat) 78.59±1.31 +4.72±1.19 -5.27±1.73 +6.27±1.15
Memory (Loc) 42.22±1.61 +26.89±1.51 +17.07±1.96 +47.39±1.00
@@
CVR-Loc-H 29.33±5.13 +17.33±5.61 +7.17±6.61 +32.00±5.48
Considering the well-established reasoning abilities of LLMs, we used self-captioning and pre-captioning methods to investigate the source of poor visual reasoning in VLMs through vision-text decoupling. The results for these experiments are displayed in Table 2; for a fully granular look, please see Appendix Table A.4.2.
@@
Notably, the largest performance gains by far for the SC method compared to the baseline were for location tasks. Across both PAM and CVR results, applying SC to location-only tasks saw an average accuracy improvement of 18.31% compared to 3.54% for category tasks. This stark difference suggests that poor VLM spatial reasoning may stem from inadequate CoT prompting. Two key observations support this: (1) Qwen2.5-VL-7B can accurately describe object locations and use those descriptions to answer location-based Perception questions, and (2) Qwen2.5-VL-7B fails at these same questions when no location information is included in the prompt. These simple observations point to a core limitation in current CoT strategies, even for single-image spatial reasoning, and likely explain the near-chance performance on multi-image CVR and PAM location tasks. By contrast, it is unclear why this bottleneck affects category tasks less. A likely reason is the abundance of object recognition data in VLM training.
To assess the broader applicability of our self-captioning (SC) method for improving VLM performance, we conducted additional experiments on GPT-4o and the largest variant of Qwen2.5-VL. The results, presented in Table A.5.1, demonstrate that the effectiveness of the SC method extends beyond Qwen2.5-VL-7B. These findings further confirm our CoT diagnosis, showing that comparable improvements can be gained in larger and SOTA proprietary VLMs.
Finally, we used the SC-I method on Qwen2.5-VL-7B to test the effect of including images in the model’s input on its performance. In this setup, the model is provided with its self-caption descriptions interleaved between the corresponding images. This method saw significant improvements on location-only PAM tasks and almost all CVR tasks. However, on category-only PAM tasks, the addition of images decreased performance relative to the baseline. Furthermore, SC-I improvements were significantly smaller than those of the image-free SC method. These results suggest the inclusion of images alone leads to diminished reasoning abilities, likely resulting from attention capacity or interference issues. These results align with recent mechanistic interpretability findings from [28] that show attention limitations are a contributing factor to spatial reasoning weaknesses in VLMs.
@@ 4.3.2 Fine-tuning VLMs
[…]
@@
Perception (Cat) 83.67±4.18 +2.00±3.97 +2.67±3.89 +6.00±3.46
Perception (Loc) 44.33±5.59 +17.00±5.48 +26.33±5.13 +34.33±4.62
Feature Attention 55.85±2.65 +20.74±2.26 +21.41±2.23 +16.96±2.37
Spatial Attention 57.48±2.63 +24.59±2.05 +22.52±2.13 +17.93±2.30
Memory (Cat) 78.59±1.31 +2.35±1.26 -3.20±1.38 +3.68±1.22
Memory (Loc) 42.22±1.61 +10.78±1.63 +22.72±1.56 +27.69±1.50
@@
CVR-Cat-H 39.33±5.49 +9.33±5.62 +26.67±5.33 +28.67±5.25
CVR-Loc-H 29.33±5.13 +9.67±5.49 +36.67±5.33 +44.33±4.96
We investigated whether fine-tuning Qwen2.5-VL-7B on random CVR tasks leads to improvements on the core cognitive tasks of the PAM dataset (Perception, Attention, and Memory; Table 3). We used supervised fine-tuning with LoRA on randomly generated CVR tasks (please see Section A.3 for method details). We found that fine-tuning on even 1,000 trials yields large gains in performance on core cognitive tasks, and the performance gains further increase with additional data. Similar to our method of self-captioning, fine-tuning appears to disproportionately improve location-based PAM task performance. Unsurprisingly, training on 100,000 held-out CVR task trials improved CVR task performance the most.
While these results are promising, standard LoRA methods are known to be prone to overfitting [58, 59, 60]. To make sure no severe overfitting has occurred in the Qwen2.5-VL-7B model after fine-tuning, we benchmark the LoRA fine-tuned models along with the base model on MMBench and MMMU-Pro. Table A.3.2 shows that supervised LoRA fine-tuning on CVR tasks can provide modest improvements to MMBench. However, performance on MMMU-Pro slightly decreases when fine-tuned on smaller dataset sizes.
@@ 5 Discussion
Fine-tuning VLMs on a small dataset of random visual reasoning tasks also led to substantial improvements along all cognitive axes while yielding slight gains on out-of-distribution benchmarks. We believe it is likely that more sophisticated methods of fine-tuning, such as Reinforcement Learning with CoT reward [61], could lead to improved generalization. This hypothesis is supported by high correlations when comparing model PAM and CVR accuracy against MMMU-Pro accuracy, which can be found in Figures A.6.1. This is an important finding that further validates our evaluation datasets and approach.
@@ 6 Limitations
Our study has several limitations that future work can address: 1) The stimulus set used to generate test samples consisted of synthetic objects from only eight categories of everyday items. Future evaluations could incorporate a wider variety of objects, including more diverse and natural visual categories. 2) Visual frames in our benchmarks featured uniformly colored backgrounds, consistent with practices in cognitive science and neuroscience. However, future studies could adopt more naturalistic settings where objects are embedded in complex scenes. Leveraging generative image models may offer a promising avenue for creating such stimuli [62, 63, 64]. 3) For feasibility, we restricted our fine-tuning experiments to the Qwen2.5-VL model and used up to 100k trials. Expanding these experiments to larger models and datasets could yield additional insights. Generative task environments, such as those proposed by Lei et al. [22], provide a convenient framework for scaling data to arbitrarily large sizes. 4) Our fine-tuning experiments consisted of tasks with a random level of complexity. While fine-tuning on these tasks already resulted in improvements on the cognitive benchmarks, we expect that a more balanced data distribution would be even more helpful. Future studies could perform fine-tuning on more complexity-balanced task datasets.
@@ A.4 Granular Results & Details
[…]
@@
[…]
Task Base SC SC-I PC
Mem-Cat-R 76.53±3.03 80.00±2.86 69.60±4.02 85.87±2.49
Mem-Cat-C 82.13±2.74 91.07±2.05 76.40±3.71 93.87±1.73
Mem-Loc-R 33.20±3.36 72.00±3.21 63.40±4.21 97.47±1.15
Mem-Loc-C 52.13±3.57 69.73±3.28 59.20±4.29 96.00±1.42
Perc-Cat-R 74.67±6.90 76.67±6.72 65.00±9.19 80.67±6.29
Perc-Cat-C 92.67±4.25 93.33±4.09 90.00±5.96 82.67±6.04
Perc-Loc-R 33.33±7.46 73.33±7.01 58.00±9.50 99.33±1.78
Perc-Loc-C 55.33±7.86 72.67±7.07 51.00±9.61 99.33±1.78
CVR-Cat-H 39.33±5.49 45.33±5.60 46.00±6.84 52.33±5.62
CVR-Loc-H 29.33±5.13 46.67±5.61 36.50±6.61 61.33±5.48
CVR-Cat-M 54.00±7.88 57.33±7.82 61.00±9.39 66.67±7.46
CVR-Loc-M 49.33±7.90 59.33±7.77 52.00±9.61 60.00±7.75
CVR-Cat-L 60.00±7.75 62.67±7.65 61.00±9.39 61.33±7.70
CVR-Loc-L 56.00±7.85 64.67±7.56 56.00±9.55 68.67±7.34
Att-Feat-R 58.67±4.53 80.44±3.66 46.00±5.60 99.78±0.60
Att-Feat-C 54.44±3.25 76.56±2.76 58.17±3.93 67.22±3.06
Att-Spa-R 49.33±4.60 72.22±4.13 40.67±5.52 83.78±3.40
Att-Spa-C 61.56±3.17 77.33±2.73 63.33±3.84 57.11±3.23
Mem-Dis-Cat-R 78.56±2.68 79.56±2.63 66.00±3.78 75.44±2.81
Mem-Dis-Cat-C 77.78±2.22 83.33±1.99 78.56±2.68 85.56±1.88
Mem-Dis-Loc-R 34.11±3.09 75.56±2.80 65.67±3.79 94.44±1.51
Mem-Dis-Loc-C 51.41±2.66 70.81±2.42 59.89±3.20 89.93±1.61

### v2

@@ Abstract
Vision-Language Models (VLMs) have shown remarkable progress in visual understanding in recent years. Yet, they still lag behind human capabilities in specific visual tasks such as counting or relational reasoning. To understand the underlying limitations, we adopt methodologies from cognitive science, analyzing VLM performance along core cognitive axes: Perception, Attention, and Memory. Using a suite of tasks targeting these abilities, we evaluate state-of-the-art VLMs, including GPT-4o. Our analysis reveals distinct cognitive profiles: while advanced models approach ceiling performance on some tasks (e.g. category identification), a significant gap persists, particularly in tasks requiring spatial understanding or selective attention. Investigating the source of these failures and potential methods for improvement, we employ a vision-text decoupling analysis, finding that models struggling with direct visual reasoning show marked improvement when reasoning over their own generated text captions. These experiments reveal a strong need for improved VLM Chain-of-Thought (CoT) abilities, even in models that consistently exceed human performance. Furthermore, we demonstrate the potential of targeted fine-tuning on composite visual reasoning tasks and show that fine-tuning smaller VLMs moderately improves core cognitive abilities. While this improvement does not translate to large enhancements on challenging, out-of-distribution benchmarks, we show broadly that VLM performance on our datasets strongly correlates with performance on established benchmarks like MMMU-Pro and VQAv2. Our work provides a detailed analysis of VLM cognitive strengths and weaknesses and identifies key bottlenecks in simultaneous perception and reasoning while also providing an effective and simple solution.
@@ 2 Related Works
Beyond these broad vision-language benchmarks, recent studies have focused on testing VLMs on more narrow cognitive domains. BlindTest [48] measures the abilities of VLMs to solve visual geometric tasks. VCog-Bench [49] takes inspiration from early neuro-developmental cognitive tests and evaluates VLMs on their ability to solve abstract logic puzzles and pattern recognition tasks. VisFactor from [50] measures similar highly abstract visual reasoning abilities. Schulze Buschoff et al. [51] focus their cognitive evaluations on intuitive physics, causal reasoning, and intuitive psychology. Visual abductive reasoning is another related area of research[52, 53]. It focuses on the inference of a hidden cause for a give observation, which differs from our focusing on logical reasoning. However, by far the most commonly measured cognitive domain is spatial reasoning [28, 29, 30, 23, 31].
@@ 3.1 Data Source
We utilize the iWISDM task environment [22] to generate all cognitive tasks and fine-tuning data in this study. This environment enables the procedural generation of an effectively limitless number of vision-language decision-making tasks. In this study, we leverage iWISDM to generate tasks with varying levels of complexity, aligning with the requirements of each cognitive axis in the PAM dataset. These tasks range from simple single-object localization to more complex ones requiring logical reasoning and object comparisons across image sequences. The PAM dataset consists of tasks manually designed to isolate core cognitive abilities. The CVR dataset was generated using iWISDM to produce complex, composite reasoning problems. Together they provide a wholistic set of tasks that allow us to systematically analyzing the cognitive profiles of state-of-the-art VLMs.
@@ 3.4 Decoupling Vision & Text via Captioning
Example prompts and Python code for our captioning methods can be found in Appendix A.11.
@@ 4.1 Cognitive Evaluation of VLMs with PAM Dataset
Table 1: Average PAM and CVR scores across all tasks for various open-source models, closed-source models, and human performance. Scores are presented as mean percentage accuracy ($\pm$ standard deviation). Best performance for open-source models for each task is bolded. Best performance overall for each task is underlined.
@@
In addition to task accuracy, we recorded human response times to provide context on the perceived difficulty of these tasks (see Appendix Table A.2.1). The response times clearly correlate with task complexity; for instance, the average time for a high-complexity category CVR task (77.21s) was more than ten times longer than for a simple perception task (7.39s). Interestingly, humans consistently took longer to solve category-based tasks than their location-based counterparts, a pattern that contrasts with the performance of most VLMs, which struggle more with spatial reasoning. A difference likely reflecting the abundance of object category based task training data in comparison to spatial task data.
@@ 4.2.1 Relationship Between PAM & CVR
Intuitively, models with stronger memory, attention, and perception abilities should also perform better on tasks that require a combination of these skills. Since our CVR tasks are designed to engage exactly these abilities in combination, we plot each model’s PAM performance against its CVR performance, as shown in Figure 2. All performances on all three core cognitive axes showed strong correlation to CVR task performance. Memory task performance stands out with a correlation of 0.94. This is likely due to the abundance of Memory comparison-based subtasks (i.e comparing two objects with irrelevant images in between), contained within CVR tasks. Furthermore, models’ PAM and CVR accuracies are significantly correlated to their performance on widely used benchmarks such as MMMU-Pro (Figures A.7.1), which further validates the effectiveness of the evaluations presented here.
@@ 4.3.1 Decoupling Vision & Text via Captioning
Table 2: PAM and CVR performance of Qwen2.5-VL-7B using different captioning methods. The Base column shows absolute percentage accuracies ($\pm$ standard deviation). Subsequent columns show the percentage change in accuracy from the Base method ($\pm$ standard deviation). Positive changes are shown in green, and negative changes in red. Asterisks denote statistical significance: * p < 0.05, ** p < 0.01, *** p < 0.001.
@@
Perception (Loc) 44.33±5.59 +28.67*±5.00 +10.17±6.84 +55.00*±1.11
Feature Attention 55.85±2.65 +22.00***±2.21 -1.74±3.25 +22.22***±2.21
Spatial Attention 57.48±2.63 +18.15***±2.29 -1.70±3.24 +8.52±2.52
Memory (Cat) 78.59±1.31 +4.72***±1.19 -5.27***±1.73 +6.27***±1.15
Memory (Loc) 42.22±1.61 +26.89***±1.51 +17.07***±1.96 +47.39***±1.00
@@
CVR-Loc-H 29.33±5.13 +17.33*±5.61 +7.17±6.61 +32.00*±5.48
Considering the well-established reasoning abilities of LLMs, we used self-captioning and pre-captioning methods to investigate the source of poor visual reasoning in VLMs through vision-text decoupling. The results for these experiments are displayed in Table 2; for a fully granular look, please see Appendix Table A.4.2. We also evaluated one-shot prompting, however this approach did not yield a significant performance boost compared to other methods, a finding that aligns well with previous literature [60, 61].
@@
Notably, the largest performance gains by far for the SC method compared to the baseline were for location tasks. Across both PAM and CVR results, applying SC to location-only tasks saw an average accuracy improvement of 18.31% compared to 3.54% for category tasks. This stark difference suggests that the bottleneck of VLM spatial reasoning may stem from inadequate CoT training. Two key observations support this: (1) Qwen2.5-VL-7B can accurately describe object locations from visual inputs, and use those descriptions to answer location-based Perception questions, and (2) Qwen2.5-VL-7B fails at the same questions when no text-based location information is provided. These simple observations point to a core limitation in current CoT strategies, even for single-image spatial reasoning, and likely explain the near-chance performance on multi-image CVR and PAM location tasks. By contrast, it is unclear why this bottleneck affects category tasks less. A likely reason is the abundance of object recognition data in VLM training.
To assess the broader applicability of our self-captioning (SC) method for improving VLM performance, we conducted additional experiments on GPT-4o and the largest variant of Qwen2.5-VL. The results, presented in Table A.5.1 and A.5.2, demonstrate that the effectiveness of the SC method extends beyond Qwen2.5-VL-7B. These findings further confirm our CoT diagnosis, showing that comparable improvements can be gained in larger and SOTA proprietary VLMs.
Finally, we used the SC-I method on Qwen2.5-VL-7B to test the effect of including images in the model’s input on its performance. In this setup, the model is provided with its self-caption descriptions interleaved between the corresponding images. This method saw significant improvements on location-only PAM tasks and almost all CVR tasks. However, on category-only PAM tasks, the addition of images decreased performance relative to the baseline. Furthermore, SC-I improvements were significantly smaller than those of the image-free SC method. These results suggest the inclusion of images alone leads to diminished reasoning abilities, likely resulting from attention capacity (see Attention Analysis in Appendix A.9) or interference issues. These results align with recent mechanistic interpretability findings from [28] that show attention limitations are a contributing factor to spatial reasoning weaknesses in VLMs.
@@ 4.3.2 Fine-tuning VLMs
Table 3: PAM and CVR performance of Qwen2.5-VL-7B base model and fine-tuned variants using LoRA with different amounts of training data. The Base column shows absolute percentage accuracies ($\pm$ standard deviation). Subsequent columns show the percentage change in accuracy from the Base method ($\pm$ standard deviation). Positive changes are shown in green, and negative in red. Asterisks denote statistical significance: * p < 0.05, ** p < 0.01, *** p < 0.001.
@@
Perception (Cat) 83.67±4.18 +2.00±3.97 +2.67±3.89 +6.00*±3.46
Perception (Loc) 44.33±5.59 +17.00±5.48 +26.33±5.13 +34.33*±4.62
Feature Attention 55.85±2.65 +20.74***±2.26 +21.41***±2.23 +16.96***±2.37
Spatial Attention 57.48±2.63 +24.59***±2.05 +22.52***±2.13 +17.93***±2.30
Memory (Cat) 78.59±1.31 +2.35**±1.26 -3.20*±1.38 +3.68**±1.22
Memory (Loc) 42.22±1.61 +10.78***±1.63 +22.72***±1.56 +27.69***±1.50
@@
CVR-Cat-H 39.33±5.49 +9.33*±5.62 +26.67*±5.33 +28.67*±5.25
CVR-Loc-H 29.33±5.13 +9.67±5.49 +36.67*±5.33 +44.33*±4.96
We investigated whether fine-tuning Qwen2.5-VL-7B on random CVR tasks leads to improvements on the core cognitive tasks of the PAM dataset (Perception, Attention, and Memory; Table 3). To prove the generalization of the fine-tuning, we also performed LoRA fine-tuning on Qwen2.5-VL-32B model on the same data (see Table A.6.1). We used supervised fine-tuning with LoRA on randomly generated CVR tasks (please see Section A.3 for method details). We found that fine-tuning on even 1,000 trials yields large gains in performance on core cognitive tasks, and the performance gains further increase with additional data. Similar to our method of self-captioning, fine-tuning appears to disproportionately improve location-based PAM task performance. Unsurprisingly, training on 100,000 held-out CVR task trials improved CVR task performance the most.
While these results are promising, standard LoRA methods are known to be prone to overfitting [62, 63, 64]. To mitigate this risk and make sure no severe overfitting has occurred in the Qwen2.5-VL-7B model during fine-tuning, we experimented with several dropout rates. Our training runs with lower dropout ratios (0.0 and 0.1) experienced overfitting, resulting in a rise in the validation loss during training. Consequently, we selected a dropout rate of 0.2 for our final models. To further assess generalization, we benchmark the LoRA fine-tuned models along with the base model on MMBench, MMMU-Pro and VQAv2. Table A.3.2, A.3.3 shows that supervised LoRA fine-tuning on CVR tasks can provide modest improvements to MMBench and VQAv2. However, performance on MMMU-Pro slightly decreases when fine-tuned on smaller dataset sizes.
@@ 5 Discussion
Fine-tuning VLMs on a small dataset of random visual reasoning tasks also led to substantial improvements along all cognitive axes while yielding slight gains on out-of-distribution benchmarks. We believe it is likely that more sophisticated methods of fine-tuning, such as Reinforcement Learning with CoT reward [65], could lead to improved generalization. This hypothesis is supported by high correlations when comparing model PAM and CVR accuracy against established benchmarks like MMMU-Pro and VQAv2, which can be found in Figures A.7.1 and A.8.1. This is an important finding that further validates our evaluation datasets and approach.
@@ 6 Limitations
Our study has several limitations that future work should address: 1) The stimulus set used to generate test samples consisted of synthetic objects from only eight categories of everyday items. Future evaluations could incorporate a wider variety of objects, including more diverse and natural visual categories. 2) Visual frames in our benchmarks featured uniformly colored backgrounds, consistent with practices in cognitive science and neuroscience. However, future studies could adopt more naturalistic settings where objects are embedded in complex scenes. Leveraging generative image models may offer a promising avenue for creating such stimuli [66, 67, 68]. 3) For feasibility, we restricted our fine-tuning experiments to the Qwen2.5-VL model and used up to 100k trials. Expanding these experiments to larger models and datasets could yield additional insights. Generative task environments, such as those proposed by Lei et al. [22], provide a convenient framework for scaling data to arbitrarily large sizes. 4) Our fine-tuning experiments consisted of tasks with a random level of complexity. While fine-tuning on these tasks already resulted in improvements on the cognitive benchmarks, we expect that a more balanced data distribution would be even more helpful. Future studies could perform fine-tuning on more complexity-balanced task datasets. 5) The self-captioning (SC) method, while effective at improving performance, introduces a computational trade-off. Our analysis shows that this method has roughly double the inference runtime compared to the standard direct-input approach, which could be a critical factor in real-world applications. A breakdown of this analysis is provided in Appendix A.10.
Acknowledgments
This research was supported by the Healthy-Brains-Healthy-Lives startup supplement grant, the NSERC Discovery grant RGPIN-2021-03035, and CIHR Project Grant PJT-191957. P.B. was supported by FRQ-S Research Scholars Junior 1 grant 310924, and the William Dawson Scholar award. M.P. was supported by the UNIQUE Masters and PhD Fellowships, and the Stichting Formation Award. All analyses were executed using resources provided by the Digital Research Alliance of Canada (Compute Canada) and funding from Canada Foundation for Innovation project number 42730. Z.W. was supported by China Scholarship Council Grant 202506070003. L.G was supported by the UNIQUE MSc Excellence Scholarship 2024-2025.
@@ A.2 Human Baseline Details
Table A.2.1: Average human response time across all task types.
Task Mean Response Time (s)
Perception (Cat) 7.39±1.56
Perception (Loc) 5.77±1.12
Feature Attention 11.34±1.78
Spatial Attention 10.79±1.89
Memory (Cat) 9.67±1.30
Memory (Loc) 8.15±1.05
CVR-Cat-L 30.68±6.47
CVR-Loc-L 23.39±6.83
CVR-Cat-M 43.81±17.41
CVR-Loc-M 34.36±9.74
CVR-Cat-H 77.21±25.90
CVR-Loc-H 41.44±10.84
@@ A.3 Supervised VLM Fine-tuning on Composite Visual Reasoning
Table A.3.3: Benchmarking scores for Qwen2.5-VL-7B and Qwen2.5-VL-72B on the VQAv2 dataset with different caption techniques and finetuning configurations.
7B Models 72B Models
Base SC LoRA 1k LoRA 10k LoRA 100k Base SC
Accuracy (%) 60.07 65.07 63.96 64.44 65.84 71.09 72.02
@@ A.4 Granular Results & Details
Table A.4.1: Average granular accuracy (%) comparison across all tasks and models. Scores are presented as mean percentage accuracy ($\pm$ standard deviation). LLaVa: LLaVa-OneVision-7B; MiniCPM: MiniCPM-V-2.6-8B; InternVL: InternVL2.5-8B; Qwen 7B: Qwen2.5-VL-7B; Qwen 72B: Qwen2.5-VL-72B; 4o-Mini: GPT-4o-Mini; 4o: GPT-4o; human: Human subjects.
@@
Table A.4.2: Average granular accuracy (%) comparison of Qwen2.5-VL-7B using different captioning methods. Scores are presented as mean percentage accuracy ($\pm$ standard deviation).
Task Base SC SC-I PC One Shot
Mem-Cat-R 76.53±3.03 80.00±2.86 69.60±4.02 85.87±2.49 78.53±2.93
Mem-Cat-C 82.13±2.74 91.07±2.05 76.40±3.71 93.87±1.73 86.80±2.42
Mem-Loc-R 33.20±3.36 72.00±3.21 63.40±4.21 97.47±1.15 32.27±3.34
Mem-Loc-C 52.13±3.57 69.73±3.28 59.20±4.29 96.00±1.42 50.80±3.57
Perc-Cat-R 74.67±6.90 76.67±6.72 65.00±9.19 80.67±6.29 76.67±6.72
Perc-Cat-C 92.67±4.25 93.33±4.09 90.00±5.96 82.67±6.04 90.00±4.84
Perc-Loc-R 33.33±7.46 73.33±7.01 58.00±9.50 99.33±1.78 27.33±7.07
Perc-Loc-C 55.33±7.86 72.67±7.07 51.00±9.61 99.33±1.78 44.00±7.85
CVR-Cat-H 39.33±5.49 45.33±5.60 46.00±6.84 52.33±5.62 31.33±5.22
CVR-Loc-H 29.33±5.13 46.67±5.61 36.50±6.61 61.33±5.48 30.00±5.16
CVR-Cat-M 54.00±7.88 57.33±7.82 61.00±9.39 66.67±7.46 58.00±7.80
CVR-Loc-M 49.33±7.90 59.33±7.77 52.00±9.61 60.00±7.75 44.00±7.85
CVR-Cat-L 60.00±7.75 62.67±7.65 61.00±9.39 61.33±7.70 61.33±7.70
CVR-Loc-L 56.00±7.85 64.67±7.56 56.00±9.55 68.67±7.34 58.00±7.80
Att-Feat-R 58.67±4.53 80.44±3.66 46.00±5.60 99.78±0.60 58.89±4.53
Att-Feat-C 54.44±3.25 76.56±2.76 58.17±3.93 67.22±3.06 54.00±3.25
Att-Spa-R 49.33±4.60 72.22±4.13 40.67±5.52 83.78±3.40 51.11±4.60
Att-Spa-C 61.56±3.17 77.33±2.73 63.33±3.84 57.11±3.23 61.89±3.17
Mem-Dis-Cat-R 78.56±2.68 79.56±2.63 66.00±3.78 75.44±2.81 77.67±2.72
Mem-Dis-Cat-C 77.78±2.22 83.33±1.99 78.56±2.68 85.56±1.88 70.89±2.42
Mem-Dis-Loc-R 34.11±3.09 75.56±2.80 65.67±3.79 94.44±1.51 29.56±2.98
Mem-Dis-Loc-C 51.41±2.66 70.81±2.42 59.89±3.20 89.93±1.61 52.22±2.66
@@
Table A.4.4: The iWISDM AutoTask parameters used to generate the PAM tasks. No and/or operators, boolean operators and switch operators are allowed in these tasks.
Complexity # of trial frames # of distractors root operators
Perception Category Report 1 0 GetCategory
Perception Category Compare 2 0 IsSame
Perception Location Report 1 0 GetLoc
Perception Location Compare 2 0 IsSame
Feature Attention Report 1 1-3 GetLoc
Feature Attention Compare 2 1-7 IsSame
Spatial Attention Report 1 1-3 GetCategory
Spatial Attention Compare 2 1-7 IsSame
Memory Category Report 2-6 0 GetCategory
Memory Category Compare 3-7 0 IsSame
Memory Location Report 2-6 0 GetLoc
Memory Location Compare 3-7 0 IsSame
Memory with Distractors Category Report 2-7 1-6 GetCategory
Memory with Distractors Category Compare 3-11 1-9 IsSame
Memory with Distractors Location Report 2-7 1-6 GetLoc
Memory with Distractors Location Compare 3-11 1-9 IsSame